OpenStudy (anonymous):

how do i find the solution for -2x-2y=4 and 8y=-8x-16. Can anybody please help me?!

OpenStudy (anonymous):

Graph them! The first equation needs to be in y = mx + b form. Then use something like geogebra to graph. The place they intersect is the solution.

OpenStudy (anonymous):

If you graph them the lines should be on top of each other. This is a result of the two equations being sort of different algebraic names for the same line. Since they both describe the same line, all coordinates (x,y) that satisfy \[-2x-2y=4\]aka\[8y=-8x-16\]aka\[x+y=-2\] make up the solution. In other words, any and all equations for the line are solutions.

OpenStudy (anonymous):

Depends if you need to solve it for x or y. If for x then isolate it on one side. If for y isolate it on one side. If you graph it you need it in either slope-intercept form or standard form.